OpenStudy (anonymous):

What is the coordinates of 7pi/4

Parth (parthkohli):

You mean the rectangular coordinates?

OpenStudy (anonymous):

\[{7\pi\over4}=\left({\sqrt{2}\over2},-{\sqrt{2}\over2}\right)\]